What is yhe answer equation to prove you are not a bot: (2+8)-4+3
Answer by lenny460(1073)

(2 + 8) - 4 + 3
(2 + 8) = 2 + 8 = 10
10 - 4 + 3
10 - 4 = 6
6 + 3 = 9
The Answer:
9


obuong3@aol.com